OpenStudy (anonymous):

fan medal and testimony help!!!. What is the solution to the equation 1.2m − 0.8 = −2.0m? m = 0.25 m = 0.4 m = 1 m = 4

Mehek (mehek14):

first add 1.2 + 2

OpenStudy (anonymous):

3.2

Mehek (mehek14):

and add 0 + 0.8

OpenStudy (anonymous):

0.8

Mehek (mehek14):

so you have \(3.2m = 0.8\) now divide \(\dfrac{0.8}{3.2}\)

OpenStudy (anonymous):

0.25

Mehek (mehek14):

\(m=0.25\)

OpenStudy (anonymous):

thanks can you help with a couple more

Mehek (mehek14):

sure

OpenStudy (anonymous):

thanks your great!

Mehek (mehek14):

^_^

OpenStudy (anonymous):

Which statement is true about the equation fraction 3 over 4z − fraction 1 over 4z + 1 = fraction 2 over 4z + 1? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (anonymous):

DO YOU KNOW HOW TO

OpenStudy (anonymous):

oops cap lock sorry

Mehek (mehek14):

\(\dfrac{3}{4z}-\dfrac{1}{4z+1}=\dfrac{2}{4z+1}\)

Mehek (mehek14):

\(\dfrac{3}{4z}=\dfrac{3}{4z+1}\)

OpenStudy (anonymous):

i dont know how to do that :(

Mehek (mehek14):

\(\dfrac{4}{3}*\dfrac{3}{4z+1}\)

OpenStudy (anonymous):

what is z

OpenStudy (anonymous):

im sorry i am a very slow learner and horrible at math

OpenStudy (anonymous):

can you explain more?

Mehek (mehek14):

hold on

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Which statement is true about the equation fraction 3 over 4z − fraction 1 over 4z + 1 = fraction 2 over 4z + 1? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (anonymous):

lol i just wanna see it cuz the other one went up

Mehek (mehek14):

I'm getting no solution

OpenStudy (anonymous):

@Mehek14 thanks for trying

Mehek (mehek14):

np

OpenStudy (anonymous):

@Nnesha can you help

OpenStudy (anonymous):

@brittanyirvinn @beckyjeanwilson @brybry245 @brandonptv3

OpenStudy (anonymous):

it this question below Which statement is true about the equation fraction 3 over 4z − fraction 1 over 4z + 1 = fraction 2 over 4z + 1? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

Nnesha (nnesha):

\[\huge\rm \frac{ 3 }{ 4z } -\frac{ 1 }{4z+1 } = \frac{ 2 }{ 4z+1 }\] like this ?

OpenStudy (anonymous):

yeah but i dont get how to do that

Nnesha (nnesha):

first of all solve left side find common denominator

OpenStudy (anonymous):

4z + 4z

Nnesha (nnesha):

when you multiply same bases you should add their exponents so 4z times 4z= ?

OpenStudy (anonymous):

16z

Nnesha (nnesha):

common denominator is 4z(4z+1) so distribute parentheses by 4z when u multiply same bases you should add exponents \[\huge\rm x^m \times x^n = x^{m+n}\]

OpenStudy (anonymous):

wow i mind is BLOWN!!!

OpenStudy (anonymous):

i dont get it :(

OpenStudy (anonymous):

im sry

Nnesha (nnesha):

4z times 4z =(4 times 4)(z times z)

Nnesha (nnesha):

z times z equal what ?

OpenStudy (anonymous):

16 and 2z

OpenStudy (anonymous):

z=8?

Nnesha (nnesha):

nope add their exponents z is same as z to the one power \[z^1 \times z^1 = z^{1+1}\]

OpenStudy (anonymous):

ok i dont get it im sorry

OpenStudy (anonymous):

if you dont want to help me thats okay cuz this will take a long time

Nnesha (nnesha):

here are examples \[\huge\rm x^2 \times x^2 = x^{2+2}\] \[\huge\rm 2^3 \times 2^3 = 2^{3+3}\] this is an exponent rule when we multiply same bases we add their exponents

OpenStudy (anonymous):

ok

Nnesha (nnesha):